OpenStudy (anonymous):

Which number equals sqrt(162) ? 3 sqrt(9) 9 sqrt(2) 9 sqrt(3) 3 sqrt(2)

OpenStudy (anonymous):

\[\sqrt{162} = \sqrt{81 \times 2} = 9\sqrt{2}\]

OpenStudy (anonymous):

cept 2 sqrt 9 isnt a choice

OpenStudy (wikiemol):

first find its prime factorizations sqrt(2*81) which is the same thing as sqrt(2)*sqrt(81)

OpenStudy (anonymous):

just realized that

OpenStudy (anonymous):

derp.

OpenStudy (anonymous):

oh sorry read that backwards

OpenStudy (anonymous):

nah, i deleted.I believe you can put it either way. i had it 2sqrt(81)

OpenStudy (anonymous):

how do you get the 81 and 2?

OpenStudy (wikiemol):

and sorry not prime factorizations, just start to factor it.

OpenStudy (anonymous):

well you divide 162 by 2. and fine 2x81= 162

OpenStudy (anonymous):

oooh thanks!

OpenStudy (wikiemol):

well you know 162 is divisible by 2, so just divide the number by 2. You just want to keep factoring until you find a square, and if you dont find a square then the square root cannot be rewritten.

OpenStudy (anonymous):

k thx